OpenStudy (anonymous):

simplify 1-(-1(1-sin^2thita)everything ova 1-sin^2thita

OpenStudy (anonymous):

any idea Sir?

OpenStudy (icecube):

(cosA+sinA)^2 +1 OR 2(1-sinAcosA)

Directrix (directrix):

Then, 1-(-1(1-sin^2thita)everything ova 1-sin^2thita can be thought of as: [1 - ( -1 ( 1- sin ² Θ ) )] /[ ( 1 - sin ² Θ)] = [1 - ( -1 (cos²Θ) ] / cos²Θ = (1 + cos²Θ) / cos²Θ = 1/cos²Θ + cos²Θ / cos²Θ = sec ² Θ + 1

OpenStudy (ash2326):

we have \[ \frac{1-(-(1-\sin^2 \theta))}{1-\sin^2\theta}\] we know that \[1-\sin^2\theta=\cos^2\theta\] so we get \[ \frac{1-(-\cos^2\theta)}{\cos^2\theta}\] we have now \[ \frac{1+\cos^2\theta}{\cos^2\theta}\] we get \[\sec^2 \theta +1\]